Use the Pythagorean Theorem to find the length of the diagonal of a piece of paper that measures 16 inches by 24 inches.
Answer by nerdybill(7384)

Use the Pythagorean Theorem to find the length of the diagonal of a piece of paper that measures 16 inches by 24 inches.
.
Pythagorean:
a^2 + b^2 = c^2
where
c is the hypotenuse
and a and b are the other two sides of the right triangle
.
16^2 + 24^2 = c^2
256 + 576 = c^2
832 = c^2
sqrt(832) = c
sqrt(4*4*2*2*13) = c
4*2*sqrt(13) = c
8*sqrt(13) inches = c (exact answer)
or approximately:
28.84 inches = c